What is 4n/2 sequence
Answer by erica65404(394)

n stands for the number of the sequence.
if you plug in 1 for n you will get the first number in the sequence.
4*1/2 = 2
the first number is 2.
4*2/2=4
the second number is 4.
4*3/2=6
the third number is 6.
and so forth.
so far we have 2,4,6... if we continue for the 4th, 5th, 6th and 7th number we will get 2,4,6,8,10,12,14...
--------------------------------------------------------------------------------
Email me at ericahigley@yahoo.com for more free step by step help.